OpenStudy (anonymous):

If f(x)=3x+5 and g(x)=2x-9 find f(x) - g(x)

OpenStudy (anonymous):

f(x) - g(x) = (3x + 5) - (2x - 9) =3x + 5 - 2x + 9 = x + 14.